OpenStudy (anonymous):

72 is what percent of 480? Write and solve an equation to solve the problem

OpenStudy (anonymous):

480*x/100=72; x=15

OpenStudy (anonymous):

I think its 15 percet do 72/480 * x/100 yo times 72 by 100 then divide by to get x

OpenStudy (anonymous):

72/480=15/100

OpenStudy (anonymous):

\[\frac{72}{480}*100=15\% \]